# Why aphids are not pests in cacao? An approach based on a predator-prey model with aging

### v1

@@ Abstract
We studied a mean-field predator-prey model with aging to simulate the interaction between aphids (Toxoptera aurantii) and syrphid larvae in cacao farms in Ilheus, Bahia. Based on the classical predator-prey model, we propose a system of differential equations with three rate equations. Unlike the original Lotka-Volterra model, our model includes two aphid population classes: juveniles (non-breeding) and adult females (asexually breeding). We obtained steady-state solutions for juvenile and adult populations by analyzing the stability of the fixed points as a function of model parameters. The results show that the absorbing state (zero prey population) is always possible, but not consistently stable. A nonzero stationary solution is achievable with appropriate parameter values. Using phase diagrams, we analyzed the stationary solution, providing a comprehensive understanding of the dynamics involved. Simulations on complete graphs yielded results closely matching the differential equations. We also performed simulations on random networks to highlight the influence of network topology on system behavior.
@@
aphids , cacao cultivation , population dynamics , mean-field approximation , differential equations
@@ 5 Conclusions
Our mean-field model introduced a system of differential equations that incorporated key biological factors, such as the maturation process of aphids and their asexual reproduction. We derived steady-state solutions and analyzed the stability of the system’s fixed points. The results showed that the absorbing state, where the aphid population is zero, is a possible solution but not always stable. A nonzero stationary solution can be obtained with an appropriate set of parameters, as confirmed by the phase diagrams.
The simulations on complete graphs validated our mean-field predictions, demonstrating that the model accurately captures the system’s dynamics under well-mixed conditions. However, simulations on random networks revealed additional complexities. These simulations exhibited variability in the exact population levels at equilibrium relative to mean-field predictions, highlighting the influence of network topology on the dynamics. Specifically, random networks can sustain larger fluctuations, indicating that local interactions and stochastic effects play a significant role in population stability.
Our findings have practical implications for pest management in agricultural settings. The presence of syrphid larvae effectively controls aphid populations, preventing them from reaching pest status. The effectiveness of this biological control strategy may vary based on the spatial distribution and connectivity of the habitat. Enhancing habitat connectivity and promoting a well-mixed predator-prey environment could improve the stability and effectiveness of natural pest control mechanisms.
Overall, this study reinforces the theoretical foundations of predator-prey dynamics while uncovering the additional layers of complexity introduced by network structures. These insights are crucial for developing more nuanced and effective strategies for ecological management and pest control in agricultural ecosystems. By understanding the mechanisms that regulate aphid populations in cacao farms, we can better design sustainable agricultural practices that leverage natural predation for pest management [5].
Future research should explore the impact of different network topologies on predator-prey dynamics to better understand the role of spatial structure and connectivity in population stability. Empirical studies in real-world agricultural settings are essential to validate our theoretical and simulation results. Additionally, investigating interactions between multiple predator and prey species and incorporating environmental factors such as climate change and agricultural practices could provide a more comprehensive understanding of ecological dynamics [56]. Understanding the spatial dynamics of populations is crucial for effective pest management. Hanski (1999) and Levins (1969) have been pioneers in the study of metapopulations and their implications for ecology and conservation [57, 58].
@@ Acknowledgements
V.R.V.A., N.G.F.M., E.N.S. and A.T.C.S. thank the Department of Education of the government of the State of Bahia for the incentive bonus for scientific production. E.N.S. thanks CAPES Foundation-Brazil for a scholarship and grants that allowed data collection in the field (BEX 1609/00-9). V.R.V.A. thanks Professor Dagoberto da Silva Freitas for encouraging him to return to scientific research. AC thanks the Centro de Ciências Agrárias of UFSCar for its support of scientific research.

### v2

@@ Abstract
We studied a mean-field predator-prey model with aging to simulate the interaction between aphids (Toxoptera aurantii) and syrphid larvae in cacao farms in Ilheus, Bahia. Based on the classical predator-prey model, we propose a system of differential equations with three rate equations. Unlike the original Lotka-Volterra model, our model includes two aphid population classes: juveniles (non-breeding) and adult females (asexually breeding). We obtained steady-state solutions for juvenile and adult populations by analyzing the stability of the fixed points as a function of model parameters. The results show that the absorbing state (zero prey population) is always possible, but not consistently stable. A nonzero stationary solution is achievable with appropriate parameter values. Using phase diagrams, we analyzed the stationary solution, providing a comprehensive understanding of the dynamics involved. Simulations on complete graphs yielded results closely matching the differential equations. We also performed simulations on random networks to highlight the influence of network topology on system behavior. Our findings highlight the critical role of life-stage structure, predation, and spatial variation in stabilizing predator-prey systems. This emphasizes the importance of network effects in population dynamics and refines the framework for biological pest control in agriculture. Ultimately, our research contributes to sustainable agricultural practices.
@@
aphids , cacao cultivation , population dynamics , mean-field approximation
@@ 5 Conclusions
Our model extends classical predator-prey frameworks by incorporating key biological factors such as the maturation process of aphids and their asexual reproduction, offering a stage-structured perspective. By distinguishing between juvenile and adult aphids, we capture more nuanced population dynamics, addressing gaps in traditional models that lack life-stage differentiation [11, 41, 42, 43, 47].
We derived steady-state solutions and analyzed the stability of the system’s fixed points. The results showed that while the absorbing state (zero aphid population) is possible, it is not always stable. A nonzero stationary solution emerges with an appropriate set of parameters, confirmed through phase diagrams.
Simulations on complete graphs validated our mean-field predictions, demonstrating that the model accurately represents the system’s dynamics under well-mixed conditions. However, simulations on random networks revealed additional complexities, such as larger fluctuations in population levels at equilibrium, which were not fully predicted by the mean-field approach. These results highlight the significant role that network topology and stochastic effects play in population stability, particularly in pest management contexts [53, 56].
Our findings contribute to theoretical biology by showing that the interplay between life-stage structure, predation, and spatial heterogeneity is crucial for understanding how predator-prey systems remain stable. These results underscore the importance of network effects in population dynamics and provide a more refined theoretical framework for biological pest control in agricultural ecosystems.
From a practical perspective, the study reinforces the effectiveness of natural biological control mechanisms. The presence of syrphid larvae controls aphid populations, preventing them from reaching pest status. Enhancing habitat connectivity and promoting a well-mixed predator-prey environment could further improve the stability and effectiveness of pest control strategies [25, 57].
Future research should explore the influence of different network topologies on predator-prey dynamics and how spatial structure impacts population stability. Empirical studies are needed to validate these theoretical and simulation results in real-world agricultural settings. Additionally, further work should investigate how interactions between multiple predator and prey species, as well as environmental factors such as climate change, affect these dynamics [58]. In conclusion, this study integrates concepts from predator-prey dynamics, network theory, and biological control, contributing to a deeper understanding of ecological interactions in agroecosystems and offering insights that can guide sustainable pest management practices [5].
@@ Acknowledgements
V.R.V.A., N.G.F.M., E.N.S. and A.T.C.S. thank the Department of Education of the government of the State of Bahia for the incentive bonus for scientific production. E.N.S. thanks CAPES Foundation-Brazil for a scholarship and grants that allowed data collection in the field (BEX 1609/00-9). V.R.V.A. thanks Professor Dagoberto da Silva Freitas for encouraging him to return to scientific research. A.C. thanks the Centro de Ciências Agrárias of UFSCar for its support of scientific research.